R3 R1 R4 iz R2 Vs ww

Vs=10,R1=Rs=8,R2=3,R3=4,R4=RL=2

Perform source transformation on Vs and R1. compute the equivalent current source and parallel resistance, as well as the voltage across and current through RL      
Answer #1

Consider the circuit:- R=42 R = 8.2 %3D Rs =82 R=R=2 Vs =lo v Any voliays Souree cam be Conver ted into curreut Sourco with a

10 . I= 8+8 = 0.625 A. value of paralle resistamco - Ŕ= 8+8 = 162. . New crcuit is B R3=42 IS IF Tst YIL Rq = R=Q2 0-625(T 1

The load current I is calculated y using curent division vule at nocle A I = i, +i2 i, = Ix R 16 = 0.625 x %3D 16+ = 0.556 A

Again at node B– I = i,xR, Re+ R,+ Rq O:556 X 3 3+4+2 IL= 0185 A. Valtnge across K, V= ILX Rq = 0-185X2 = 0:37 V.

Please like and rate the answer.

If you have any doubts regarding answer feel free to ask in the comment section.